OpenStudy (anonymous):

Can someone help me with this problem? Two displacement vectors when added produce a resultant vector of magnitude 100 km pointing 37 degrees south of east. One displacement vector is 200 km in a direction 60 degrees north of east. Determine the magnitude and direction of the other displacement vector. How can i find the other displacement vector? :3 can someone help me?

OpenStudy (festinger):

 magnitude can be found using \[\sqrt{x^2+y^2}=c\] while angle can be found using \[tan \theta = \frac{y}{x}\]